# USTC-TD: A Test Dataset and Benchmark for Image and Video Coding in 2020s

### v2

@@ Abstract
Image/video coding has been a remarkable research area for both academia and industry for many years. Testing datasets, especially high-quality image/video datasets are desirable for the justified evaluation of coding-related research, practical applications, and standardization activities. We put forward a test dataset namely USTC-TD, which has been successfully adopted in the practical end-to-end image/video coding challenge of the IEEE International Conference on Visual Communications and lmage Processing (VCIP) in 2022 and 2023. USTC-TD contains 40 images at 4K spatial resolution and 10 video sequences at 1080p spatial resolution, featuring various content due to the diverse environmental factors (e.g. scene type, texture, motion, view) and the designed imaging factors (e.g. illumination, lens, shadow). We quantitatively evaluate USTC-TD on different image/video features (spatial, temporal, color, lightness), and compare it with the previous image/video test datasets, which verifies the wider coverage and more diversity of the proposed dataset. We also evaluate both classic standardized and recent learned image/video coding schemes on USTC-TD with PSNR and MS-SSIM, and provide an extensive benchmark for the evaluated schemes. Based on the characteristics and specific design of the proposed test dataset, we analyze the benchmark performance and shed light on the future research and development of image/video coding. All the data are released online: https://esakak.github.io/USTC-TD .
@@ I Introduction
End-to-end image/video compression has been a research focus on visual data compression for both academia and industry for over six years[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. A number of technologies have been developed, such as expressive auto-encoder neural networks, precise probability estimation neural networks, and conditional end-to-end coding frameworks, and so on. Until recently, the performances of both end-to-end image and video compression schemes have surpassed that of the advanced H.266/VVC under certain test conditions [10, 11, 26, 27].
For the evaluation of these image/video compression schemes in the practical application and standardization, they are usually benchmarked with objective and subjective metrics to evaluate their rate-distortion (RD) performance and a trade-off between coding efficiency and reconstructed quality. To sufficiently consider the effectiveness of these quality assessments, the test results of representative image/video test datasets are the key to reflecting the practicability and generalization of the researcher’s scheme.
In this paper, a new image/video dataset, named USTC-TD, is proposed for testing and evaluating the practical image/video coding algorithms. USTC-TD contains 40 images and 10 video sequences with a wide content coverage. For the image dataset, each image has a high spatial resolution (4K) in RGB, YUV444/420 color space, and PNG/YUV file format. For the video dataset, each video sequence consists of 96 frames, and each frame is captured at 30 frames per second (fps) with 1080p spatial resolution in RGB, YUV444/420 color space, and PNG/YUV file format. For the construction of datasets, the data is collected with the specific design of different content factors (environmental/imaging-related factors), which aims to cover as close as possible to the real-world coding transmission scenes. Compared with the previous common test image/video datasets [28, 29, 30, 31, 32, 33, 34], we use different quantitative criteria to comprehensively evaluate the diversity of the proposed dataset from the perspective of spatial, temporal, colorfulness, and lightness information, and demonstrate the wide coverage for various image/video features.
@@
In addition, we establish baselines and evaluate the classic standardized compression schemes [35, 1, 2, 3, 4, 5, 36] and recent learned image/video compression schemes [6, 7, 8, 9, 10, 11, 12, 16, 19, 20, 21, 22, 23, 25, 26, 27] under different metrics (PSNR, MS-SSIM[37]), and then benchmark and analyze their performance on the proposed dataset to shed light on the future research and development of image/video coding. The benchmark data and test scripts are online available with the proposed dataset and released on the open-sourced website for researchers to reproduce conveniently.
@@
  • We conduct a comprehensive evaluation of the proposed dataset by using different quantitative criteria, and demonstrate the wide coverage of various image/video features.
  • We conduct a comprehensive evaluation of the advanced image/video compression schemes on the proposed dataset, and establish an extended baseline for the evaluated image/video coding schemes benchmarked on USTC-TD.
@@
The remainder of the paper is structured as follows. Section II mentions the background of previous compression-related test datasets. Section III summarises the data collection process of the proposed dataset. Section IV introduces the construction of the image/video dataset of USTC-TD, and analyzes the characteristics of the proposed dataset. Section V presents the experimental configuration and the evaluation of the advanced compression schemes, and further analyzes their performance, limitation, and inspiration. Section VI concludes the paper and presents some suggestions for future work. Finally, we mention the copyright for the usage of USTC-TD.
@@ II Background
Image Compression Test Dataset. For the evaluation of previous image compression schemes, Kodak[28], Tecnick[29] (sampling setting), CLIC (professional setting) [30] are commonly used, the setting is mentioned in the Table I , and the introduction is summarised below:
@@
  • CLIC[29] is a high-quality image set collected from Unsplash[38], and contains images of similar quality from potentially different sources. It has been successfully applied in the workshop and challenge on learned image compression (CLIC) of IEEE/CVF Computer Vision and Pattern Recognition Conference (CVPR) and Data Compression Conference (DCC).
@@
  • UVG[31] contains 16 test video sequences. They are captured with Sony F65 video camera in 16-bit F65RAW-HFR format and converted to YUV420 videos by ffmpeg tool[39]. It is widely used in the evaluation of advanced learned video compression methods[12, 16, 19, 20, 21, 22, 23, 25, 26, 27].
@@ III-A Camera and Format Configuration
The images and video sequences are captured by using Nikon-D3200 and Nikon-Z-fc for USTC-TD 2022 and 2023 datasets, and the specific camera parameters are shown in Table III and Table IV . For the format of images and video sequences in the dataset, they are transcoded from Raw camera format (DNG, MOV) and then converted to RGB, YUV 4:4:4, YUV 4:2:0 color space/format by using the ffmpeg [39] tool and the specific conversion standard of color space (BT.601, BT.709 [26, 40]).
@@ III-B Collection Configuration
To develop a comprehensive and diverse image/video dataset, we consider the various content factors of collected data, including the environmental factors (e.g. scene type, texture, motion, view), imaging factors (e.g. illumination, lens, shadow), which cover as close as possible to the real-world coding transmission scenes. For each factor, we categorize it into different types implicit in the construction of our dataset, the categories are shown in Table V . According to these conditions, we choose more than twenty different scene types (e.g. dormitory, library, river bank, institutes, parks, classroom, street, vehicles), and adjust different camera parameters to capture. For each image, we take ten shots of the same scene at the same time to select the better one. For each video, we record five minutes for each scene with the same range of joint sense to select the ninety-six frames.
@@
IV Dataset Construction and Analysis
@@ IV-A Construction of USTC-TD 2022 and 2023 Image Dataset
Based on the characteristics of previous image datasets[28, 29, 30], our proposed dataset aims to cover various scenarios. Considering the various content factors, we combine different environmental factors and imaging factors in the collection process. For the diversity of environmental factors, we consider the scene type, texture, and view factors. For the diversity of imaging factors, we consider the resolution, illumination, and shadow factors. In Fig.  1 , we show the all collected image data of USTD-TD 2022, and in Table VI , we show the specific configuration of each image, and make it convenient to the researchers’ scheme design for different application scenes. The collected image data and configuration of USTC-TD 2023 are also mentioned in Fig.  1  and Table VII . Based on USTD-TD 2022, USTC-TD 2023 considers more extreme factors in real-world scenes.
@@
USTC-2022-06 Scenery 4096×2160 Geometric Appropriate Exposure Upward Level Cast
@@
USTC-2022-13 Scenery 4096×2160 Structural Appropriate Exposure Upward Level Cast
USTC-2022-14 People, Vehicle 4096×2160 Natural Appropriate Exposure Horizontal Level Cast
@@
USTC-2022-18 People, Building 4096×2160 Natural Appropriate Exposure Horizontal Level Soft
USTC-2022-19 People, Building 4096×2160 Natural Overexposure Horizontal Level Soft
@@
USTC-2023-01 People, Room 3840×2160 Natural Appropriate Exposure Horizontal Level Hard
USTC-2023-02 Scenery 3840×2160 Geometric Underexposure Overhead Level Cast
USTC-2023-03 Scenery 3840×2160 Structural Underexposure Horizontal Level Soft
@@
USTC-2023-19 Close Shot 3840×2160 Natural Appropriate Exposure Horizontal Level Hard
@@
Figure 2: Illustration of the frame 01, 06, 11 of each video sequence in video test dataset of USTC-TD 2023.
@@
USTC-BasketballDrill YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-BasketballPass YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-BicycleDriving YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural Horizontal Level Moving
@@
USTC-ParkWalking YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural Horizontal Level Moving
USTC-Running YUV420, 444, RGB Complex People, Sport 1920×1080 High Natural Horizontal Level Moving
USTC-ShankingHand YUV420, 444, RGB Complex People, Daliy life 1920×1080 High Natural Horizontal Level Moving
@@ IV-B Construction of USTC-TD 2023 Video Dataset
In Fig. 2 , we show the partial frames of all collected video sequences in the USTD-TD 2023 dataset, and the specific configuration of each video in Table VIII , which make it convenient for the scheme design of research in real-world scenes under different temporal correlation types. Note that the different motion types are obviously classified in the Table VIII with different colors.
@@
Different from the previous video dataset, we add more specific temporal correlation types in our proposed video dataset. For example, in USTC-BicycleDriving, we capture the video with a fast scene change (lens motion), high-speed moving objects, and object occlusion, which is a challenge for many inter-alignment schemes of submitted solutions in VCIP Challenge 2023 [48] and many optical flow-based video compression schemes [12, 49, 16, 19, 20, 21, 22, 23, 25, 26, 27] (the detailed performance analysis is mentioned in Section V.B). For the performance of different schemes of this sequence, the learned video compression schemes are far inferior to the traditional codecs [1, 2, 3, 4, 5]. In USTC-Snooker, we capture the scenes with the tiny motion and fast lens motion, which is also a challenge for optical flow-based schemes. At present, most of the learned video compression schemes use the optical flow-based alignment[50, 51], the optical flow-based motion estimation is difficult to capture the tiny motion and further influences their performance. Therefore, we put forward our proposed video dataset with the above specific designs, and hope the efficient testing datasets can help the researchers discover the problem related to temporal characteristics on their video compression scheme.
@@
Figure 3: The visualization of the evaluation of spatial information (SI) and colorfulness (CF) features on different image test datasets. Scatter diagram represents the SI versus CF, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of images under different SI scores.
@@
Figure 4: The visualization of the evaluation of lightness information (LI) and CF features on different image test datasets. Scatter diagram represents the LI versus CF, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of images under different LI scores.
@@
Figure 5: The visualization of the evaluation of temporal information (TI) and SI features on different video test datasets. Scatter diagram represents the TI versus SI, and corresponding convex hulls indicates the coverage of different datasets. The histogram represents the number of videos under different TI scores.
IV-C Analysis of USTC-TD Dataset

### v3

@@ Abstract
Image/video coding has been a remarkable research area for both academia and industry for many years. Testing datasets, especially high-quality image/video datasets are desirable for the justified evaluation of coding-related research, practical applications, and standardization activities. We put forward a test dataset namely USTC-TD, which has been successfully adopted in the practical end-to-end image/video coding challenge of the IEEE International Conference on Visual Communications and lmage Processing (VCIP) in 2022 and 2023. USTC-TD contains 40 images at 4K spatial resolution and 10 video sequences at 1080p spatial resolution, featuring various content due to the diverse environmental factors (e.g. scene type, texture, motion, view) and the designed imaging factors (e.g. illumination, lens, shadow). We quantitatively evaluate USTC-TD on different image/video features (spatial, temporal, color, lightness), and compare it with the previous image/video test datasets, which verifies its excellent compensation for the shortcomings of existing datasets. We also evaluate both classic standardized and recently learned image/video coding schemes on USTC-TD using objective quality metrics (PSNR, MS-SSIM, VMAF) and subjective quality metric (MOS), providing an extensive benchmark for these evaluated schemes. Based on the characteristics and specific design of the proposed test dataset, we analyze the benchmark performance and shed light on the future research and development of image/video coding. All the data are released online: https://esakak.github.io/USTC-TD .
@@ I Introduction
End-to-end image/video compression has been a research focus on visual data compression for both academia and industry for over six years[6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. A number of technologies have been developed, such as expressive auto-encoder neural networks, precise probability estimation neural networks, and conditional end-to-end coding frameworks, and so on. Until recently, the performances of both end-to-end image and video compression schemes have surpassed that of the advanced H.266/VVC under certain test conditions [10, 11, 26, 27].
For the evaluation of these image/video compression schemes in the practical application and standardization, they are usually benchmarked with objective and subjective quality metrics to evaluate their rate-distortion (RD) performance and a trade-off between coding efficiency and reconstructed quality. To sufficiently consider the effectiveness of these quality assessments, the test results of representative image/video test datasets are the key to reflecting the practicability and generalization of the researcher’s scheme.
In this paper, a new image/video dataset, named USTC-TD, is proposed for testing and evaluating the practical image/video coding algorithms. USTC-TD contains 40 images and 10 video sequences with a wide content coverage. For the image dataset, each image is captured with a high spatial resolution (4K) and converted into RGB, YUV444/420 color space, and PNG/YUV file format. For the video dataset, each video sequence consists of 96/300 frames, and each frame is captured at 30 frames per second (fps) with 1080p spatial resolution and converted into RGB, YUV444/420 color space, and PNG/YUV file format. For the construction of image and video datasets, the data is collected with the specific design of different content factors (environmental/imaging-related factors), which aims to cover as close as possible to the real-world coding transmission scenes. Compared with the common test image/video datasets [32, 33, 34, 35, 36, 37, 38], we use different quantitative criteria to comprehensively evaluate the diversity of the proposed USTC-TD from the perspective of spatial, temporal, colorfulness, and lightness information, demonstrating its excellent coverage and effectively compensating for the shortcomings of existing datasets.
@@
In addition, we establish baselines and evaluate the classic standardized compression schemes [1, 2, 3, 4, 5] and recently learned image/video compression schemes [6, 7, 8, 9, 10, 11, 12, 16, 19, 20, 21, 22, 23, 25, 26, 27, 28, 29, 30, 31] under objective quality metrics (PSNR, MS-SSIM[39], VMAF [^1] Available online at  https://github.com/Netflix/vmaf.[40]) and subjective quality metric (MOS) [41, 42, 43, 44]), and then benchmark and analyze their performance on the proposed dataset to shed light on future research and development of image/video coding. The benchmark data and test scripts are available online with the proposed dataset and released on the open-sourced website for researchers to reproduce conveniently.
@@
  • We conduct a comprehensive evaluation of the proposed dataset by using different quantitative criteria, demonstrating the excellent compensation of USTC-TD for existing image/video datasets.
  • We conduct a comprehensive evaluation of the advanced image/video compression schemes on the proposed dataset, and establish an extended baseline for the evaluative image/video coding schemes benchmarked on USTC-TD.
@@
The remainder of the paper is structured as follows. Section II mentions the background of previous compression-related test datasets. Section III summarises the data collection process of the proposed dataset. Section IV introduces the construction of the image and video dataset of USTC-TD, and discusses the characteristics and utilization of the proposed dataset. Section V presents the experimental configuration and the evaluation of the advanced compression schemes on USTC-TD, and further analyzes their performance, limitation, and inspiration. Section VI concludes the paper and presents some suggestions for future work.
@@ II Background
Image Compression Test Dataset. For the evaluation of previous image compression schemes, Kodak [^2] Available online at  https://r0k.us/graphics/kodak/.[32], Tecnick[33] (sampling setting), CLIC (professional setting) [34] are commonly used, the setting is mentioned in Table I , and the introduction is summarised below:
@@
  • CLIC[34] is a high-quality image set collected from Unsplash [^3] Available online at  https://unsplash.com/., and contains images of similar quality from potentially different sources. It has been successfully applied in the workshop and challenge on learned image compression (CLIC) of IEEE/CVF Computer Vision and Pattern Recognition Conference (CVPR) and Data Compression Conference (DCC).
@@
  • UVG[35] contains 16 test video sequences. They are captured with Sony F65 video camera in 16-bit F65RAW-HFR format and converted to YUV420 videos by ffmpeg tool  [^4] Available online at  https://ffmpeg.org/.. It is widely used in the evaluation of advanced learned video compression methods[12, 16, 19, 20, 21, 22, 23, 25, 26, 27, 28, 29, 30, 31].
@@ III-A Camera and Format Configuration
The images and video sequences are captured by using Nikon-D3200 and Nikon-Z-fc for USTC-TD 2022 and 2023 datasets, and the specific camera parameters are shown in Table III and Table IV . For the format of images and video sequences in the dataset, they are transcoded from Raw camera format (DNG, MOV) and then converted to RGB, YUV444/420 color space/format by using the ffmpeg tool and the conversion standard of color space (BT.601[45]).
@@ III-B Collection Configuration
To develop a comprehensive and diverse image/video dataset, we consider the various content factors of collected data, including the environmental factors (e.g. scene type, texture, motion, view), imaging factors (e.g. illumination, lens, shadow), which cover as close as possible to the real-world coding transmission scenes. For each factor, we categorize it into different types implicit in the construction of our dataset, the categories are shown in Table V . According to these conditions, we choose more than twenty different scene types (e.g. dormitory, library, river bank, institutes, parks, classroom, street, vehicles), and adjust different camera parameters to capture. For each image, we take ten shots of the same scene at the same time to select the best one. For each video, we record five minutes for each scene with the same range of joint sense to select the short and long test sequences.
@@
IV Dataset Construction, Analysis, Discussion
@@ IV-A Construction of USTC-TD 2022 and 2023 Image Dataset
Based on the characteristics of previous image datasets[32, 33, 34], our proposed dataset aims to cover various scenarios. Considering the various content factors, we combine different environmental factors and imaging factors in the collection process. For the diversity of environmental factors, we consider the scene type, texture, and view factors. For the diversity of imaging factors, we consider the resolution, illumination, and shadow factors. In Fig.  1 , we show all collected image data of USTD-TD 2022, and in Table VIII , we show the specific configuration of each image, and make it convenient for the researchers’ scheme design for different application scenes. The collected image data and configuration of USTC-TD 2023 are also mentioned in Fig.  1  and Table VIII . Based on USTD-TD 2022, USTC-TD 2023 considers more extreme factors in real-world scenes.
@@
      USTC-2022-06       Scenery       4096×2160       Natural, Geometric       Appropriate Exposure       Upward Level       Cast
@@
      USTC-2022-13       Scenery       4096×2160       Nature, Structural       Appropriate Exposure       Upward Level       Cast
      USTC-2022-14       People, Vehicle       4096×2160       Natural       Underexposure       Horizontal Level       Cast
@@
      USTC-2022-18       People, Building       4096×2160       Natural, Geometric       Appropriate Exposure       Horizontal Level       Soft
      USTC-2022-19       People, Building       4096×2160       Natural, Geometric       Overexposure       Horizontal Level       Soft
@@
     USTC-2023-01      People, Room      3840×2160      Natural, Structural      Appropriate Exposure      Horizontal Level      Hard
     USTC-2023-02      Scenery      3840×2160      Geometric, Structural      Underexposure      Overhead Level      Cast
     USTC-2023-03      Scenery      3840×2160      Natural, Structural      Underexposure      Horizontal Level      Soft
@@
     USTC-2023-19      Close Shot      3840×2160      Natural      Underexposure      Horizontal Level      Hard
@@
USTC-BasketballDrill YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural, Geometric Horizontal Level Moving
USTC-BasketballPass YUV420, 444, RGB Medium People, Sport 1920×1080 High Natural, Geometric Horizontal Level Moving
USTC-BicycleDriving YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural, Structural Horizontal Level Moving
@@
USTC-ParkWalking YUV420, 444, RGB Complex People, Daily Life 1920×1080 High Natural, Structural Horizontal Level Moving
USTC-Running YUV420, 444, RGB Complex People, Sport 1920×1080 High Natural, Structural Horizontal Level Moving
USTC-ShakingHands YUV420, 444, RGB Complex People, Daliy life 1920×1080 High Natural, Geometric Horizontal Level Moving
@@
Figure 2: Illustration of each video sequence in USTC-TD 2023 video dataset. The 0$\sim$96/0$\sim$300 frames correspond to the short and long setting.
@@ IV-B Construction of USTC-TD 2023 Video Dataset
In Fig. 2 , we show the partial frames of all collected video sequences in the USTD-TD 2023 dataset, and the specific configuration of each video in Table VIII , which make it convenient for the scheme design of research in real-world scenes under different temporal correlation types. Note that the different motion types are obviously classified in the Table VIII  with different colors. For the construction of USTC-TD video dataset, two kinds of settings are constructed: short setting and long setting. For the short setting (3 seconds), a 96-frame subset is selected from each full-length captured test video sequence (frame rate: 30fps, 300 seconds, 9000 frames in total) to reduce the high complexity associated with long sequences during the practical testing, which can promote the fast evaluation in the research process. For the long setting, we extend from the last frame of the short setting and directly enlarge the short sequence to 300 frames (10 seconds), which aligns with the 10-second setting of test sequence length of HEVC and VVC common test condition (CTC).
@@
Different from the previous video dataset, we add more specific temporal correlation types in our proposed video dataset. For example, in USTC-BicycleDriving, we capture the video with a fast scene change (lens motion), high-speed moving objects, and object occlusion, which is a challenge for many inter-alignment schemes of submitted solutions in VCIP Challenge 2023  [^5] Available online at  https://vcip2023.iforum.biz/page/goto/. and many optical flow-based video compression schemes [12, 53, 16, 19, 20, 21, 22, 23, 25, 26, 27] (the detailed performance analysis is mentioned in Section V.B). For the performance of different schemes of this sequence, the learned video compression schemes are far inferior to the traditional codecs [1, 2, 3, 4, 5]. In USTC-Snooker, we capture the scenes with the tiny motion and fast lens motion, which is also a challenge for optical flow-based schemes. At present, most of the learned video compression schemes use the optical flow-based alignment[54, 55], and the optical flow-based motion estimation is difficult to capture the tiny motion and further influences their performance. Therefore, we put forward our proposed video dataset with the above specific designs, and hope the efficient testing datasets can help the researchers discover the problem related to temporal characteristics on their video compression scheme.
@@
Figure 3: The visualization of the evaluation of spatial information (SI) and colorfulness (CF) features on different image test datasets. Scatter diagram represents the SI versus CF, and corresponding convex hulls indicate the coverage of different datasets. The histogram represents the number of images under different SI scores.
@@
Figure 4: The visualization of the evaluation of lightness information (LI) and CF features on different image test datasets. Scatter diagram represents LI versus CF, and corresponding convex hulls indicate the coverage of different datasets. The histogram represents the number of images under different LI scores.
@@
Figure 5: The visualization of the evaluation of temporal information (TI) and SI features on different video test datasets. Scatter diagram represents the TI versus SI, and corresponding convex hulls indicate the coverage of different datasets. The histogram represents the number of videos under different TI scores.
IV-C Analysis of USTC-TD Image and Video datasets
In the above subsections, the construction of each dataset of USTC-TD is introduced, here we point out the specific characteristics of the proposed dataset and highlight the driving factors behind its benefit, and discuss its practical application.
IV-C1 Characteristics of USTC-TD Image and Video datasets